OpenStudy (anonymous):

Factor the polynomial. X^3 -3x^2+x-3

OpenStudy (anonymous):

2

OpenStudy (mertsj):

Group the terms. Factor x^2 out of the first two terms

OpenStudy (anonymous):

May you factor it for me please

OpenStudy (mertsj):

\[(x^3-3x^2)+(x-3)=x^2(x-3)+1(x-3)=(x-3)(x^2+1)\]

OpenStudy (anonymous):

\[x^3 -3x^2+x-3\] \[(x^3 -3x^2)+(x-3)\] \[x^2(x -3)+ 1(x-3)\] \[(x^2 + 1)(x-3)\]

OpenStudy (anonymous):

Well thanks a lot i appreciated

OpenStudy (anonymous):

yw :)